# W-RAG: Weakly Supervised Dense Retrieval in RAG for Open-domain Question Answering

### v1

@@ 3.2.1. DPR
DPR utilizes a bi-encoder architecture, where the question and passage are independently mapped to an embedding space through two separate BERT encoders. For a given question $q$, the relevance score $R_{q,s}$ for each passage $s$ is given by the similarity between their respective [CLS] token embeddings, $\mathbf{e}_{q}$ and $\mathbf{e}_{s}$. We use cosine similarity as the similarity function:
@@
Following the in-batch negative training method introduced in DPR, we process each ranked list $\mathcal{S}\in\mathcal{W}$ by extracting the top-ranked passage $s_{i}$ along with its associated question $q_{i}$. These pairs are then grouped into batches of size $n$ to form training batches, optimizing the retriever using a Multiple Negatives Ranking (MNR) loss (Henderson et al., 2017). In this approach, for the $i$th question-passage pair, $s_{i}$ is treated as the positive example, while the other passages within the same batch are treated as negative examples. The loss function is thus:
@@
The scaler $\alpha$ is used to amplify the cosine similarity score, usually set at 20 according to the default setting in sentence-transformers  [^2] https://sbert.net/docs/package_reference/sentence_transformer/losses.html#multiplenegativesrankingloss.
@@ 3.2.2. ColBERT
ColBERT employes a bi-encoder architecture with late interaction, where the question and passage are independently encoded using a shared BERT model to obtain embeddings $E_{q}$ and $E_{s}$. The relevance score for each passage is computed on the question side using a “MaxSim” operation, which sums the maximum similarities between any passage token embedding with each question token embedding:
(5) $R_{q,s}=\sum_{i\in|E_{q}|}\max_{j\in|E_{s}|}\text{dot}(E_{q_{i}},E_{s_{j}})$
Here, $\text{dot}(x,y)$ represents the dot product between $x$ and $y$. The training dataset for ColBERT consists of triplet samples $\langle q,s^{+},s^{-}\rangle$, extracted from each ranked list $\mathcal{S}$. In these triplets, $s_{r_{1}}$ is treated as the positive passage $s^{+}$, while $s_{r_{2}},s_{r_{3}},\dots,s_{r_{m+1}}$ are treated as negative passages $s^{-}$. Here, $m$ represents the number of hard negatives selected from $\mathcal{S}$. ColBERT produces two scores $R_{q,s^{+}}$ and $R_{q,s^{-}}$ from each triplet, and the model is optimized through a pairwise softmax cross-entropy loss:
@@ 4.1. Task and Datasets
To assess the effectiveness of W-RAG, we study three crucial components: the quality of the LLM-generated weak labels, the retrieval performance of the weakly trained dense retriever, and the overall effectiveness of the RAG system on the OpenQA task. We conduct experiments on four well-known datasets for OpenQA: MSMARCO QnA v2.1 (Nguyen et al., 2016), NQ (Kwiatkowski et al., 2019), SQuAD (Rajpurkar et al., 2016), and WebQ (Berant et al., 2013). MSMARCO QnA v2.1 shares the same corpus as MSMARCOv1 Passage Retrieval, with questions originating from real user queries submitted to Bing, making the language conversational. NQ also features real user questions and uses a corpus of documents from the English Wikipedia. In this work, we use a chuncked version of NQ as prepared by Karpukhin et al. (Karpukhin et al., 2020a). SQuAD’s corpus is similarly derived from the English Wikipedia, where the passages were first retrieved and then sampled. The questions and answers are manually written by crowdworkers, so they do not reflect natural language as closely as questions from real user queries like MSMARCO and NQ. WebQ’s corpus is drawn from Freebase, a large knowledge graph. Due to the nature of knowledge graphs, WebQ’s questions and answers are entity-related and factoid-based, thus making them less conversational compared to other datasets like MSMARCO and NQ.
For each dataset, we uniformly random sampled 5,000 question-answer pairs and a corpus of 500,000 passages from the training set, ensuring that all questions had relevant passages included in the corpus. We then split the question-answer pairs into 2,000 as training set, 1,000 as validation set, and 2,000 as test set. This sampling was necessitated by resource and time constraints. We argue that this small training set is sufficient to demonstrate statistically significant differences in both retrieval and the final OpenQA performance. While our results are indicative, our method could benefit from further validation using larger datasets to ensure generalizability.
@@ 4.1.1. Weak Label Quality
For our main experiments, we use Llama3-8B-Instruct (Dubey and .etl, 2024) to serve as the reranker. The 2,000 question-answer pairs in the training set are used to generate weak labels. For each question, weak labels are generated by scoring the top 100 passages retrieved by BM25, based on the LLM’s likelihood of generating the answer. This process requires the LLM to perform inference 100 times to produce a ranked list. The choice of 100 passages is a trade-off between accuracy and latency. We found that BM25’s Recall@100 reaches approximately 80% across all four datasets, and retrieving additional passages yields diminishing returns as the LLM’s inference time increases linearly. We evaluate the reranking performance using different prompts and LLMs, which will be elaborated on later in the ablation studies (5.4).
@@ 4.1.2. Weakly Trained Retriever
After the LLM weakly labels the training set, we use these labels to train DPR and ColBERT. For DPR, since in-batch negatives are used during training, we only need to select the top-ranked passage to curate a list of relevant question-passage pairs, thus forming 2,000 training samples. For all four datasets, we used two different initializaiton settings for DPR: one is using bert-base-uncased to train from scratch (denoted as “DPRbase”), and the other by resuming training from an unsupervised dense retriever using Yibin-Lei/ReContriever (Lei et al., 2023b) (denoted as “DPRReCon”).
ColBERT’s training process differs slightly from that of DPR because it requires the sampling of negative passages. For each question, we select the top-ranked passage from the reranked list as the positive passage and the immediate following 10 passages as hard negatives. These hard negatives are relevant but not sufficiently informative to elicit the correct answer from the LLM. We did not explore varying the number of hard negatives, as this is not the primary focus of this work. Additionally, since the off-the-shelf ColBERTv2 is trained on the entire MSMARCO dataset, using it as the starting point would introduce data leakage. For consistency, we always train ColBERT from scratch using bert-base-uncased for all four datasets (denoted as “ColBERTbase”).
@@ 4.1.3. OpenQA Performance
Once the dense retriever is trained, we integrate it into the generic RAG pipeline, where it is used to retrieve the top-$K$ evidence passages for a given user question. The retrieved passages are directly inserted into the prompt, as shown in Figure 2 alongside the question and instructions asking the LLM to generate an answer based on the retrieved passages or, if the passages are not useful, to answer using its internal knowledge. For our main experiments, we restrict Llama3-8B-Instruct to generate a maximum of 20 tokens and only use top 1 retrieved passage to supplement the prompt. We also explored the impact of using different numbers of supplemental evidence passages on both OpenQA performance and latency, which we discuss in the ablation studies (5.4). We did not experiment with different LLMs for the OpenQA task, as the LLM component of RAG is not the primary focus of our study.
@@ 4.2. Baselines
Since our retriever is trained from scratch with limited data, we do not compare it to state-of-the-art retrievers or RAG systems. Instead, we focus on the improvement from untrained baselines to our weakly trained models and compare them with models trained on ground-truth data.
For OpenQA evaluation, we use Llama3-8B-Instruct as the LLM in the RAG pipeline, assessing the impact of different retrievers on final OpenQA performance. We compare these retrievers against the same baseline models used for OpenQA. Additionally, we introduce two baselines: “Naive,” where Llama3 answers the question without any external information, showcasing its parameter knowledge, and “Groundtruth,” where the ground-truth passage is inserted into the prompt, representing the best possible OpenQA performance.
To assess weak label quality, we compare Llama3-8B-Instruct with BM25 after Llama3 generates weak labels on the training set of each dataset. For retriever performance, we evaluate two categories of models. The first category includes entirely unsupervised models: BM25, ColBERT initialized with bert-base-uncased (denoted as “ColBERTinit”), Contriever (Izacard et al., 2021), and ReContriever (Lei et al., 2023b). Contriever learns by contrasting sampled passages from different documents, assuming that each document is somewhat unique, while ReContriever enhances this by estimating relevance among sampled passages. The second category includes DPRReCon and ColBERTbase, both trained on ground-truth data. For a fair comparison, we use the same questions and an equal number of positive and negative samples as training data.
Retriever MSMARCO QnA NQ SQuAD WebQ
F1 Rouge-L BLEU-1 F1 Rouge-L BLEU-1 F1 Rouge-L BLEU-1 F1 Rouge-L BLEU-1
Naive 0.2749 0.2392 0.2019 0.1158 0.1136 0.0618 0.0885 0.0845 0.0508 0.1512 0.1487 0.0823
Groundtruth 0.4677 0.4310 0.3541 0.2186 0.2171 0.1193 0.2691 0.2675 0.1613 0.1736 0.1707 0.0937
ColBERTinit 0.2185 0.1867 0.1619 0.0525 0.0518 0.0283 0.0501 0.0482 0.0296 0.0434 0.0423 0.0233
BM25 0.3060 0.2712 0.2292 0.1374 0.1358 0.0748 0.1387 0.1369 0.0835 0.1085 0.1055 0.0594
Contriever 0.3125 0.2798 0.2347 0.1446 0.1432 0.0796 0.1144 0.1124 0.0680 0.1273 0.1257 0.0702
ReContriever 0.3171 0.2854 0.2387 0.1440 0.1429 0.0789 0.1257 0.1233 0.0753 0.1274 0.1239 0.0695
Trained on Groundtruth Data
ColBERTbase 0.3227 0.2876 0.2421 0.1494 0.1482 0.0814 0.1524 0.1497 0.0900 0.1006 0.1001 0.0538
DPRReCon 0.3424 0.3102 0.2588 0.1641 0.1628 0.0902 0.1550 0.1531 0.0919 0.1397 0.1378 0.0771
Trained on W-RAG Data
ColBERTbase 0.3150 0.2803 0.2365 0.1313 0.1301 0.0712 0.1509 0.1489 0.0897 0.1003 0.0989 0.0545
DPRReCon 0.3397 0.3074 0.2559 0.1528 0.1515 0.0837 0.1425 0.1402 0.0850 0.1304 0.1276 0.0727
Table 1. OpenQA Performance. Top 1 passage is inserted into the OpenQA prompt, limit number of generated tokens to 20. : Performance improvements are statistically significant with a ttest of $p<0.05$. DPRReCon’s baseline is ReContriever, ColBERTbase’s baseline is ColBERTinit
@@ 4.3. Experimental Settings
We begin by retrieving the top 100 passages using BM25 with the rank_bm25  [^3] https://github.com/dorianbrown/rank_bm25 BM25Okapi with $k_{1}$=1.5, $b$=0.75 and $\epsilon$=0.25. The nltk  [^4] https://github.com/nltk/nltk tokenizer is used to tokenize the questions and passages. We use BEIR  [^5] https://github.com/beir-cellar/beir to manage data, and the DPR models are trained using sentence-transformers with the following hyperparameters: batch size of 128, learning rate of $2\times 10^{-5}$, AdamW optimizer, 20 epoch, and model weights are saved based on the best Recall@5 on the validation set. ColBERT is trained from scratch using RAGatouille  [^6] https://github.com/bclavie/RAGatouille with a batch size of 64, 10 hard negatives per positive, learning rate of $1\times 10^{-5}$, AdamW optimizer, and 1 epoch. Generating weak labels took approximately two days per dataset on an Nvidia V100 GPU. Our experiments are easily reproducible since BM25, ColBERT, DPR, Llama3, and all four datasets are publicly accessable.
The prompts used to generate weak labels are shown in Figure 2. We use the term “DOCUMENT” instead of “PASSAGE” because our testing experiments showed that “DOCUMENT” consistently yields better reranking performance. We suspect this is because “DOCUMENT” is a more common term and likely appears more frequently in Llama3’s training data, leading it to a deeper understanding of the word “DOCUMENT”. Additionally, we tested different orderings of the passage, question, and instructions, finding that the Passage-Question-Instruction format consistently outperforms Passage-Instruction-Question and Instruction-Passage-Question.

### v2

@@ 3.2.1. DPR
DPR utilizes a bi-encoder architecture, where the question and passage are independently mapped into a shared embedding space through two separate BERT encoders. In most practical implementations, a single shared BERT encoder is used for both the question and the passage, which is also the configuration used in our experiments. For a given question $q$ and passage $s$, the relevance score $R_{q,s}$ is given by the similarity between their respective [CLS] token embeddings, denoted as $\mathbf{e}_{q}\in\mathbb{R}^{d}$ and $\mathbf{e}_{s}\in\mathbb{R}^{d}$. We use cosine similarity as the similarity function:
@@
Following the in-batch negative training method introduced in DPR, we process each ranked list $\mathcal{S}\in\mathcal{W}$ by extracting the top-ranked passage $s_{i}$ along with its associated question $q_{i}$. These pairs are then grouped into batches of size $n$ to form training batches, optimizing the retriever using a Multiple Negatives Ranking (MNR) loss (Henderson et al., 2017). In this approach, for the $i$th question-passage pair, $s_{i}$ is treated as the positive example, while the other passages within the same batch $\{s_{j}\}_{j\neq i}$ are treated as negative examples. The MNR loss function is thus:
@@
The scaler $\alpha$ amplifies the cosine similarity score, typically set to 20 according to the default setting in sentence-transformers  [^2] https://sbert.net/docs/package_reference/sentence_transformer/losses.html#multiplenegativesrankingloss.
@@ 3.2.2. ColBERT
ColBERT employs a bi-encoder architecture with late interaction, where the question and passage are independently encoded using a shared BERT model to obtain embeddings $E_{q}\in\mathbb{R}^{l_{q}\times d}$ and $E_{s}\in\mathbb{R}^{l_{s}\times d}$, where $l_{q}$ and $l_{s}$ are the number of tokens in the question and passage, respectively, and $d$ is the embedding dimension. To compute the relevance score between a question and a passage, ColBERT applies a “MaxSim” operation: for each question token embedding $E_{q_{i}}\in E_{q}$, it finds the passage token embedding $E_{s_{j}}\in E_{s}$ that produces the highest dot product with $E_{q_{i}}$, and then sums these maximum values across all question tokens:
(5) $R_{q,s}=\sum_{i=1}^{l_{q}}\max_{j\in[1,l_{s}]}\text{dot}(E_{q_{i}},E_{s_{j}})$
Here, $\text{dot}(x,y)$ represents the dot product between vectors $x$ and $y$. The training dataset for ColBERT consists of triplet samples $\langle q,s^{+},s^{-}\rangle$, extracted from each ranked list $\mathcal{S}$. In each triplet, $s_{r_{1}}$ is treated as the positive passage $s^{+}$, while $s_{r_{2}},s_{r_{3}},\dots,s_{r_{m+1}}$ are treated as negative passages $s^{-}$. Here, $m$ represents the number of hard negatives selected from $\mathcal{S}$. ColBERT computes relevance scores $R_{q,s^{+}}$ and $R_{q,s^{-}}$ from each triplet, and optimizes the model using a pairwise softmax cross-entropy loss:
@@
Retriever NQ SQuAD WebQ MSMARCO QnA
F1 EM BLEU-1 F1 EM BLEU-1 F1 EM BLEU-1 F1 EM BLEU-1
Lower and upper bound of OpenQA performance
Naive 10.14 0.15 5.64 22.40 11.62 19.99 38.96 17.30 31.84 26.92 0.25 20.19
Ground truth 69.95 54.90 65.43 77.69 58.10 72.99 52.39 33.76 46.08 43.40 9.95 28.44
Without fine-tuning
BM25 52.32 38.40 48.08 47.01 32.90 43.62 36.63 20.68 31.38 30.53 5.16 19.81
Contriever 54.84 41.05 50.54 41.28 28.85 38.22 39.28 19.62 33.16 30.84 5.28 19.79
ColBERTinit 17.95 11.60 15.91 16.82 10.90 15.27 20.94 12.24 18.13 19.24 1.80 12.79
ReContriever 55.01 40.75 50.64 43.50 30.10 40.19 37.79 20.89 31.74 31.11 5.42 20.06
Fine-tuned with ground-truth data
ColBERTbase 53.95 39.85 49.65 52.43 37.37 48.77 34.21 19.20 29.35 30.51 4.60 19.79
DPRReCon 58.53 43.60 53.90 50.36 35.05 46.57 42.31 23.42 35.91 31.64 5.32 20.53
Fine-tuned with W-RAG data
ColBERTbase 50.35 37.75 46.42 52.28 36.55 48.32 32.16 18.99 27.69 29.79 4.95 19.01
DPRReCon 56.40 42.00 51.98 49.71 34.95 45.99 41.59 23.63 35.30 32.41 5.23 20.85
Table 1. OpenQA performance using Llama3.1-8B-Instruct. The top 5 passages retrieved by each retriever are inserted into the QA prompt. Bold indicates the best performance; Underlined indicates the second best (excluding “Naive” and “Ground truth”): Performance improvements are statistically significant with a t-test of $p<0.05$. Prompt can be found in Figure 3.
@@ 4.1. Task and Datasets
To assess the effectiveness of W-RAG generated weak labels, we study three crucial components: (1) the quality of the LLM-generated weak labels, (2) the retrieval performance of the weakly trained dense retriever, and (3) the end-to-end performance of the RAG system on the OpenQA task. We conduct experiments on four well-known datasets for OpenQA: MSMARCO QnA v2.1 (Nguyen et al., 2016), NQ (Kwiatkowski et al., 2019), SQuAD (Rajpurkar et al., 2016), and WebQ (Berant et al., 2013). MSMARCO QnA v2.1 shares the same corpus as MSMARCOv1 Passage Retrieval, with questions originating from real user queries submitted to Bing. The language is generally conversational, with answers that are typically longer. NQ also features real user questions and uses a corpus of documents from the English Wikipedia. The answers are often very short and clean. In this work, we use a chunked version of the NQ corpus as prepared by Karpukhin et al. (Karpukhin et al., 2020a). SQuAD’s corpus is similarly derived from the English Wikipedia, where the passages were first retrieved and then sampled. The questions and answers are manually written by crowdworkers, so they do not reflect natural language as closely as those from real user queries, such as MSMARCO QnA and NQ, and the answers are typically concise and focused. WebQ’s corpus is drawn from Freebase, a large knowledge graph. Due to the nature of knowledge graphs, WebQ’s questions and answers are entity-related and factoid-based, making them less conversational. Similar to NQ and SQuAD, the answers in WebQ are generally short.
For each dataset, we uniformly random sampled 5,000 question-answer pairs and a corpus of 500,000 passages from the training set, ensuring that all questions had relevant passages included in the corpus. We then split the question-answer pairs into 2,000 as training set, 1,000 as validation set, and 2,000 as test set. This sampling was necessitated by resource and time constraints. We argue that this small training set is sufficient to demonstrate statistically significant improvements in both retrieval and the final OpenQA performance. While our results are indicative, our method could benefit from further validation using more weakly labeled training data to ensure generalizability.
@@ 4.1.1. Weak Label Quality
For our main experiments, we use Llama3-8B-Instruct (Dubey and .etl, 2024) to serve as the weak supervision labeler. The 2,000 question-answer pairs in the training set are used to generate weak labels. For each question, we score the top 100 passages retrieved by BM25 based on the LLM’s likelihood of generating the ground-truth answer. We chose BM25 for its effectiveness and efficiency, though in principle this first-stage retrieval can be done with any retriever. The labeling process requires the LLM to perform 100 inference calls using the prompt shown in Figure 2 to produce a ranked list. The choice of 100 passages is a trade-off between accuracy and latency. We found that BM25’s Recall@100 reaches approximately 80% across all four datasets, and retrieving additional passages yields diminishing returns as the LLM’s inference time increases linearly. We evaluate the reranking performance using different LLMs, which will be elaborated on later in the ablation studies (5.4).
Question Answering Prompt [System Prompt] You are a helpful assistant who answers questions using the provided passage. If the passage does not contain the necessary information, you will answer the question directly based on your general knowledge. Only provide the final answer, no explanations. [User Prompt] Passage: {passage #1} Passage: {passage #2} Passage: {passage #3} Passage: {passage #4} Passage: {passage #5} Question: {question} [Assistant]
Figure 3. Prompt for RAG OpenQA where top 5 retrieved passages are used.
@@ 4.1.2. Weakly Trained Retriever
After the LLM weakly labels the training set, we use these labels to train DPR and ColBERT. For DPR, since we use in-batch negatives during fine-tuning, we only need to select the top-ranked passage to curate a list of relevant question-passage pairs, thus forming 2,000 training samples. For all four datasets, we experiment with two different initialization settings for DPR. The first uses bert-base-uncased as the starting point and fine-tunes with W-RAG data from the pre-trained base BERT model (denoted as “DPRbase”). The second initializes from an unsupervised pre-trained dense retriever, Yibin-Lei/ReContriever (Lei et al., 2023b), and fine-tunes it with W-RAG data (denoted as “DPRReCon”).
ColBERT’s training process differs slightly from that of DPR, as it requires sampling of negative passages. For each question, we select the top-ranked passage from the reranked list as the positive passage and the immediate following 10 passages as hard negatives. According to the definition of our weak labels, these hard negatives are relevant but not sufficiently informative to elicit the correct answer from the LLM. We did not explore varying the number of hard negatives, as this is not the primary focus of this work. Additionally, since the off-the-shelf ColBERTv2 is trained on the entire MSMARCO dataset, using it as the starting point would introduce data leakage. For consistency, we always fine-tune ColBERT from the pre-trained base BERT model bert-base-uncased for all four datasets (denoted as “ColBERTbase”).
@@ 4.1.3. OpenQA Performance
Once the dense retriever is trained, we integrate it into the generic RAG pipeline, where the retriever is used to index the evidence corpus and retrieve the top-$k$ evidence passages for a given user question. The top 5 retrieved passages are directly inserted into the prompt, as shown in Figure 3. Llama3.1-8B-Instruct is used to perform OpenQA, and we restrict the maximum output token to be 20 tokens, following (Jin et al., 2024). We also explored the impact of using different numbers of supplemental evidence passages on both OpenQA performance and latency, which we discuss in the ablation studies (5.4). We did not experiment with different LLMs for the OpenQA task, as the LLM component of RAG is not the primary focus of our study.
Figure 4. Same information as Table 1, comparing the impact of retrievers fine-tuned with weak labels versus ground-truth labels on OpenQA performance, both of which are fine-tuned from the baseline model. ReContriever serves as the baseline for the DPR-family; ColBERT initialized from base BERT serves as the baseline for the ColBERT-family.
@@ 4.2. Baselines
Since our retriever is fine-tuned with only limited data, we do not compare it to state-of-the-art retrievers or RAG systems. Instead, we focus on the improvement from starting point baselines to our weakly supervised models and compare them with models fine-tuned on ground-truth data. For OpenQA evaluation, we use Llama3.1-8B-Instruct as the LLM in the RAG pipeline, assessing the impact of the final OpenQA performance using top passages retrieved by different retrievers. Additionally, we introduce two baselines: “Naive,” where Llama3 answers the question without any external information, showcasing its parameter knowledge, and “Ground truth,” where the ground-truth passage is inserted into the prompt, representing the best possible OpenQA performance in our setup.
To assess the quality of weak labels, we compare the weakly labeled ranking lists with BM25 on the training set of each dataset. For the fine-tuned retriever’s retrieval performance, we evaluate three categories of models. The first category includes entirely unsupervised models: BM25, ColBERT initialized with bert-base-uncased (denoted as “ColBERTinit”), Contriever (Izacard et al., 2021), and ReContriever (Lei et al., 2023b). Contriever learns by contrasting sampled passages from different documents, assuming passages sampled from the same document are more similar than those sampled from different documents. ReContriever enhances Contriever by adding an estimated prior of the relevance among sampled passages. The second category includes retrievers fine-tuned on ground-truth data, and the third category are fine-tuned with W-RAG data. For a fair comparison between the second and third categories, we use the same questions and an equal number of positive and negative samples as training data.
@@ 4.3. Experimental Settings
We begin by retrieving the top 100 passages using BM25 with the rank_bm25  [^3] https://github.com/dorianbrown/rank_bm25 BM25Okapi with $k_{1}$=1.5, $b$=0.75 and $\epsilon$=0.25. The nltk  [^4] https://github.com/nltk/nltk tokenizer is used to tokenize the questions and passages. The prompt used to generate weak labels is shown in Figure 2. The weak relevance score between a question and a passage is extracted from the logits of the answer tokens, which is why the ground-truth answer is explicitly included in the prompt. We use BM25’s retrieval performance as a baseline and try different orderings of the passage, question, and instructions, finding that the weak labels generated from the “Passage-Question-Instruction” format consistently yield better retrieval metrics compared to the “Passage-Instruction-Question” and “Instruction-Passage-Question” formats. For fine-tuning, we borrow BEIR’s code  [^5] https://github.com/beir-cellar/beir to manage the training data. The DPR models are fine-tuned using sentence-transformers with the following hyperparameters: a batch size of 128, a learning rate of $2\times 10^{-5}$, the AdamW optimizer, 20 epochs, and model weights are saved based on the best Recall@5 on the validation set. ColBERT is fine-tuned from base BERT models using RAGatouille  [^6] https://github.com/bclavie/RAGatouille with a batch size of 64, 10 hard negatives per positive, learning rate of $1\times 10^{-5}$, AdamW optimizer, and one epoch. Generating weak labels took approximately two days per dataset on a single Nvidia V100 GPU. Our experiments are easily reproducible since BM25, ColBERT, DPR, Llama3-8B-Instruct, Llama3.1-8B-Instruct, and all four datasets are publicly accessible.